OpenStudy (anonymous):

The operating costs of a swimsuit distributor are $120,000 per year plus $23 per swimsuit. The current partners have 500,000 to spend and are considering a new partner who has $150,000. How many swimsuits can they produce with and without the new partner?

OpenStudy (igreen):

120,000 + 23x = 500,000 This is our equation.

OpenStudy (anonymous):

thank you :)

OpenStudy (igreen):

Let's solve for \(x\) to see how many swimsuits they can produce without the new partner. \(120,000 + 23x = 500,000\) \(23x = 380,000\) \(x \approx 16521.74\)

OpenStudy (igreen):

Well, \(x \approx 16,521\)

OpenStudy (igreen):

Our equation changes with the new partner: \(120,000 + 23x = 500,000 + 150,000\) Simplify: \(120,000 + 23x = 650,000\) Solve: \(120,000 + 23x = 650,000\) \(23x = 530,000\) \(x \approx 23,043\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @laurenbisko